Lim n-infinity (4^n + 5^n)^1/n.      Plz help me to solve this one ...

Grade:

2 Answers

Rinkoo Gupta
9 years ago
lim n->infinity (5n)^(1/n)[1+(4/5)n]1/n
=lim n->infinity 5[1+(4/5)n]1/n=5(1+0)0=5(1)0=5(1)=5
Ans. is 5
Thanks & Regards
Rinkoo Gupta
AskIITiansFaculty
Arun Kumar IIT Delhi
9 years ago
Hello
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty